OpenStudy (anonymous):

Solve: 2(x + 4) − (x + 1) = 10 3 1 2 0

OpenStudy (tyteen4a03):

Hint: Expand 2(x + 4) first.

OpenStudy (anonymous):

2x + 8 - x - 1 = 10 (2x - x) + 8 - 1 = 10 x + 7 = 10 x =10 - 7 x = 3 Ans

OpenStudy (tyteen4a03):

@Aleem123 Best to have the asker work out the steps herself :)

OpenStudy (anonymous):

@tyteen4a03 .. ok :)